OpenStudy (goformit100):

There are deer and peawingspans in a zoo. By counting heads they are 80. The number of their legs is 200. How many peawingspans are there ? A. 20 B. 30 C. 50 D. 60

OpenStudy (anonymous):

How many legs peawingspan has??

OpenStudy (dmezzullo):

@waterineyes can i put wut i think it is then u correct me?

OpenStudy (dmezzullo):

wut is a peawingspan anyways?

OpenStudy (goformit100):

its a bird

OpenStudy (dmezzullo):

Oh so it has 2 legs then

OpenStudy (anonymous):

Okay.. Yeah Sure @dmezzullo

OpenStudy (goformit100):

yes then ?

OpenStudy (dmezzullo):

Ok so theres 40 peawingspans and 40 deer, so deer have 4 legs, so 40 times 4 = 160 then the peawingspans have 2 legs and there are 40 of them so 20 so then there is ur other 40 u were missing so the answer would be 40. but where did i go wrong @waterineyes

OpenStudy (anonymous):

Just system of linear equations: x : deer and y : peawingspan.. x + y = 80 4x + 2y = 200 So: 4x + 4y = 320 4x + 2y = 200 2y = 120 y = 60..